OpenStudy (anonymous):

A car travels 100 meters in 5 seconds. What is the speed of the car? speed = distance over time; 1 mile = 1609 meters a.8.9 miles per hour b.16 miles per hour c.20 miles per hour d.44.7 miles per hour

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

v=d/t v= 100/5 convert m/s into miles/hour

OpenStudy (isaiah.feynman):

\[s=\frac{ d }{ t }\]Simple.

OpenStudy (anonymous):

is it b. cause that was my answer

OpenStudy (anonymous):

3600 seconds in an hour 1609 metres in a mile

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

i don't understand how you set it up

OpenStudy (anonymous):

you would first find the speed in m/s by v=d/t =20 m/s then use that value as m/s and convert to miles 20/1609 after you have your mile value, convert seconds to hour by 60x60 =360 v=d/t/mile x 3600

OpenStudy (anonymous):

it should be 3600 seconds in an hour**

OpenStudy (anonymous):

it says 160900 when i did it

OpenStudy (anonymous):

@bukloon

OpenStudy (anonymous):

20/1609x3600= 44.7

OpenStudy (anonymous):

oh now i see how you got it thanks for the help